# Scaling limit of the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant boundary conditions

### v2

@@ 5 Partition function in the scaling limit
where the Hamiltonian $\mathbb{H}$ is given by (1.3) with $q=\mbox{e}^{\frac{{\rm i}\pi}{n+2}}$ and $n\geq 0$, while the trace is taken over the $2^{2L}$ dimensional space of states: $\mathscr{V}_{2L}=\mathbb{C}_{1}\otimes\mathbb{C}_{2}\otimes\ldots\otimes\mathbb{C}_{2L}$. Keeping the ratio
@@
$Z^{({\rm scl})}={\rm Tr}_{\cal H}\Big{(}{\tt q}^{\hat{H}_{\rm CFT}}\Big{)}\qquad{\rm with}\qquad{\tt q}=\mbox{e}^{-2\pi\tau}\,.$ (5.4)
It involves the ‘CFT Hamiltonian’ $\hat{H}_{\rm CFT}$ which when restricted to the finite dimensional spaces ${\cal H}^{(\rm{cont},{\tt d})}_{p,s}$ or ${\cal H}^{(\rm{disc},{\tt d})}_{p,s}$ appearing in the decomposition of ${\cal H}$ coincides with the identity operator multiplied by the factor
$E_{\rm CFT}=\frac{p^{2}}{n+2}+\frac{s^{2}}{n}-\frac{1}{12}+{\tt d}\ .$ (5.5)
@@
$Z^{({\rm disc})}={\rm Tr}_{{\cal H}^{({\rm disc})}}\,\Big{(}{\tt q}^{\hat{H}_{\rm CFT}}\Big{)}\ ,\qquad\qquad Z^{({\rm cont})}={\rm Tr}_{{\cal H}^{({\rm cont})}}\,\Big{(}{\tt q}^{\hat{H}_{\rm CFT}}\Big{)}\ .$ (5.8)
@@
$\chi_{a,{\cal S}}({\tt q})={\tt q}^{-\frac{(\frac{n}{2}-{\cal S}-a)^{2}}{n}+\frac{p^{2}}{n+2}-\frac{1}{12}}\ \bigg{(}\prod_{j=1}^{\infty}(1-{\tt q}^{j})^{-2}\bigg{)}\,\sum_{m=0}^{\infty}(-1)^{m}\,\big{(}1-{\tt q}^{(1+m)(2{\cal S}+1)}\big{)}\,{\tt q}^{ma+\frac{m(m+1)}{2}}\ ,$ (5.9)
where, aside from the prefactor, the function $\chi_{a,{\cal S}}({\tt q})$ coincides with the generating function for the dimensions of the level subspaces ${\cal H}_{{\cal S},{\cal S}}^{({\rm disc,{\tt d}})}$, see eqs. (4.20) and (4.7). Then, the contribution of the discrete spectrum to the partition function reads as:
$Z^{({\rm disc})}=\sum_{{\cal S}\geq 0}(2{\cal S}+1)\bigg{(}\chi_{-{\cal S},{\cal S}}({\tt q})+2\!\!\!\!\!\sum_{a\in\mathbb{Z}\atop 0<a+{\cal S}<\frac{n}{2}}\!\!\!\!\!\chi_{a,{\cal S}}({\tt q})\bigg{)}\ .$ (5.10)
Each term in the sum over ${\cal S}$ has multiplicity $(2{\cal S}+1)$ as a result of the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ symmetry of the lattice model. Also, for every state with given $s=s_{a}$ (4.17) there exists another one with $s=-s_{a}$ which yields the same contribution to the partition function, except for the case when $s=\pm\frac{{\rm i}n}{2}$, where they are identified as the same state. This explains why the functions $\chi_{a,{\cal S}}({\tt q})$ come with a factor of two except the one with $a=-{\cal S}$ (recall that $s_{a}={\rm i}(\tfrac{n}{2}-{\cal S}-a)$ and hence $s_{a}=\frac{{\rm i}n}{2}$ for $a=-{\cal S}$).
@@
$Z^{({\rm cont})}=\sum_{{\cal S}\geq 0}(2{\cal S}+1)\,\int_{-\infty}^{\infty}\mbox{d}s\ \sum_{{\tt d}\geq 0}\rho^{({\tt d})}_{p}(s)\,{\tt q}^{\frac{s^{2}}{n}+\frac{p^{2}}{n+2}-\frac{1}{12}+{\tt d}}\,.$ (5.11)
Here $\rho^{({\tt d})}_{p}(s)$ is the density of states defined in formulae (4.5) and (4.6), while recall that $2p=2{\cal S}+1-(n+2)$. Notice that $Z^{({\rm cont})}$ becomes singular as $L\to\infty$:
@@
$Z^{({\rm sing})}=\sqrt{\frac{n}{2\tau}}\ \frac{\log\big{(}2^{\frac{n+2}{n}}\,L/L_{0}\big{)}}{\pi\,{\tt q}^{\frac{1}{24}}\prod_{m=1}^{\infty}(1-{\tt q}^{m})}\ \sum^{\infty}_{\mathcal{S}=0}\,\,(2\mathcal{S}+1)\ {\tt q}^{-\frac{1}{24}+\frac{p^{2}}{n+2}}\ \frac{1-{\tt q}^{2\mathcal{S}+1}}{\prod_{m=1}^{\infty}(1-{\tt q}^{m})}\,.$ (5.13)
@@
$\sum^{\infty}_{\mathcal{S}=0}\,\,(2\mathcal{S}+1)\ {\tt q}^{-\frac{1}{24}+\frac{p^{2}}{n+2}}\ \frac{1-{\tt q}^{2\mathcal{S}+1}}{\prod_{m=1}^{\infty}(1-{\tt q}^{m})}\,,$ (5.14)
@@ 6 Discussion
In this work the universal behaviour of the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant boundary conditions imposed was considered. We focused on the so-called self-dual case in the critical regime (1.2). The problem was reformulated in the Hamiltonian picture, where a central rôle belongs to $\mathbb{H}$ (1.3), which commutes with the transfer-matrix of the vertex model. The study of the $1/L$ corrections to the $L\to\infty$ behaviour of the energy of $\mathbb{H}$ for the low energy states allowed us to extract the spectrum of scaling dimensions of the statistical system. Our treatment, involving the use of novel numerical and analytic techniques, represents an advance on the type of analysis of the scaling limit of integrable lattice systems with open boundary conditions that typically exists in the literature.
The numerical construction of the RG trajectories was achieved via the method of the $Q$ operator. A key rôle was played by the formula (2.23) for the matrix elements of $\mathbb{Q}(\zeta)$, valid for a one parameter family of open boundary conditions in the sector $S^{z}=0$. This is an original result of our work. The advantage of (2.23), as opposed to the expressions for $\mathbb{Q}(\zeta)$ appearing in the literature [28, 29, 30, 31], is that it contains no infinite sums; works literally for any (generic) complex values of $q$ and the boundary parameter $\epsilon$; and can be programmed efficiently on the computer. We believe that in view of the important applications of the $Q$ operator, it would be worthwhile to extend (2.23) to the other sectors of the Hilbert space with $S^{z}\neq 0$ and the case of more general open boundary conditions.
@@
As was already observed in refs. [22, 23], the set of scaling dimensions of the statistical system possesses a continuous component, labeled by the quantum number $s=\operatorname*{slim}_{L\to\infty}b(L)\in\mathbb{R}$ with $b(L)$ from eq. (3.6) [23]. One of the results of this work is the explicit formula for the density of states $\rho^{({\tt d})}(s)$ (4.5), (4.6), which characterizes the continuous spectrum. In addition, we studied the RG trajectories $\{|\Psi_{L}\rangle\}$, where $s$ becomes a pure imaginary number in the scaling limit. Building on the analysis of [23], the discrete set $\Sigma\equiv\Sigma^{+}\cup\Sigma^{-}$ (4.19) of all admissible values of pure imaginary $s$ was found. We also determined the dimension of the linear span of states occuring in the scaling limit of $|\Psi_{L}\rangle$ with given $s\in\Sigma$ and conformal dimensions $\Delta$ (3.13).
Our work includes a full characterization of the linear space ${\cal H}$ appearing in the scaling limit of the space of low energy states of the lattice system. To describe it, ${\cal H}$ was decomposed into a direct sum of the ‘continuous spectrum’ ${\cal H}^{({\rm cont})}$ and the ‘discrete spectrum’ ${\cal H}^{({\rm disc})}$. The former, when expressed in terms of finite dimensional spaces involves a direct integral over $s$, see eqs. (4.23), (4.10), (4.11), while the latter contains a direct sum (4.18). We explained how the graded linear spaces $\oplus_{{\tt d}=0}^{\infty}{\cal H}_{p,s}^{({\rm cont},{\tt d})}$ and $\oplus_{{\tt d}=0}^{\infty}{\cal H}_{p,s}^{({\rm disc},{\tt d})}$ are irreps. of the ${\cal W}_{\infty}$ algebra – the algebra of extended conformal symmetry of the model.
@@ Acknowledgments
The authors are grateful to M. Flohr and V. Schomerus for their interest in the work and for valuable discussions. GK would also like to thank V. V. Bazhanov, S. L. Lukyanov and J. Teschner for useful and stimulating interactions.
@@ Appendix A The asymptotic coefficients $\mathfrak{C}_{p,s}^{(\pm)}$
Here we provide a closed form expression for the coefficients $\mathfrak{C}_{p,s}^{(\pm)}(\bm{w})$ that were obtained in the work [45]. Among other things, they enter into the quantization condition (3.40), (3.41).
@@
$\mathfrak{C}^{(\pm)}_{p,s}(\bm{w})=\mathfrak{C}_{p,s}^{(0,\pm)}\,\check{\mathfrak{C}}^{(\pm)}_{p,s}(\bm{w})\,,$ (A.1)
@@
while $\check{\mathfrak{C}}^{(\pm)}$ are one for ${\tt d}=0$. In the general case, they are given by the determinant of a ${\tt d}\times{\tt d}$ matrix:
ˇ$\check{\mathfrak{C}}^{(\pm)}_{p,s}(\bm{w})=\frac{(\mp 1)^{\tt d}\ \det\big{(}w_{a}^{b-1}\,{U}^{(\pm)}_{a}(b)\big{)}}{\prod_{a=1}^{\tt d}w_{a}\ \prod_{b>a}(w_{b}-w_{a})\ \prod_{a=1}^{\tt d}\big{(}2p+2a-1\pm 2{\rm i}s\big{)}}$ (A.3)
@@
${U}^{(\pm)}_{a}(D)$ $=$ $(D-1)^{2}-\bigg{(}2p+2+n\mp 2w_{a}+\sum_{b\not=a}^{\tt d}\frac{4w_{a}}{w_{a}-w_{b}}\bigg{)}\ (D-1)$
@@
$+$ $\bigg{(}\sum_{b\not=a}^{\tt d}\frac{2w_{a}}{w_{a}-w_{b}}\,\bigg{)}^{2}+\,\big{(}2\,(2p+1+n\mp 2w_{a})-n\,\big{)}\,\sum_{b\not=a}^{\tt d}\frac{w_{a}}{w_{a}-w_{b}}\ .$

### v3

@@ 5 Partition function in the scaling limit
where the Hamiltonian $\mathbb{H}$ is given by (1) with $q=\mbox{e}^{\frac{{\rm i}\pi}{n+2}}$ and $n\geq 0$, while the trace is taken over the $2^{2L}$ dimensional space of states: $\mathscr{V}_{2L}=\mathbb{C}^{2}_{1}\otimes\mathbb{C}^{2}_{2}\otimes\ldots\otimes\mathbb{C}^{2}_{2L}$. Keeping the ratio
@@
$Z^{({\rm scl})}={\rm Tr}_{\cal H}\Big{(}\texttt{q}^{\hat{H}_{\rm CFT}}\Big{)}\qquad{\rm with}\qquad\texttt{q}=\mbox{e}^{-2\pi\tau}\,.$ (5.4)
It involves the ‘CFT Hamiltonian’ $\hat{H}_{\rm CFT}$ which when restricted to the finite dimensional spaces ${\cal H}^{(\rm{cont},\texttt{d})}_{p,s}$ or ${\cal H}^{(\rm{disc},\texttt{d})}_{p,s}$ appearing in the decomposition of ${\cal H}$ coincides with the identity operator multiplied by the factor
$E_{\rm CFT}=\frac{p^{2}}{n+2}+\frac{s^{2}}{n}-\frac{1}{12}+\texttt{d}\ .$ (5.5)
@@
$Z^{({\rm disc})}={\rm Tr}_{{\cal H}^{({\rm disc})}}\,\Big{(}\texttt{q}^{\hat{H}_{\rm CFT}}\Big{)}\ ,\qquad\qquad Z^{({\rm cont})}={\rm Tr}_{{\cal H}^{({\rm cont})}}\,\Big{(}\texttt{q}^{\hat{H}_{\rm CFT}}\Big{)}\ .$ (5.8)
@@
$\chi_{a,{\cal S}}(\texttt{q})=\texttt{q}^{-\frac{(\frac{n}{2}-{\cal S}-a)^{2}}{n}+\frac{p^{2}}{n+2}-\frac{1}{12}}\ \bigg{(}\prod_{j=1}^{\infty}(1-\texttt{q}^{j})^{-2}\bigg{)}\,\sum_{m=0}^{\infty}(-1)^{m}\,\big{(}1-\texttt{q}^{(1+m)(2{\cal S}+1)}\big{)}\,\texttt{q}^{ma+\frac{m(m+1)}{2}}\ ,$ (5.9)
where, aside from the prefactor, the function $\chi_{a,{\cal S}}(\texttt{q})$ coincides with the generating function for the dimensions of the level subspaces ${\cal H}_{{\cal S},{\cal S}}^{({\rm disc,\texttt{d}})}$, see eqs. (4.20) and (4.7). Then, the contribution of the discrete spectrum to the partition function reads as:
$Z^{({\rm disc})}=\sum_{{\cal S}\geq 0}(2{\cal S}+1)\bigg{(}\chi_{-{\cal S},{\cal S}}(\texttt{q})+2\!\!\!\!\!\sum_{a\in\mathbb{Z}\atop 0<a+{\cal S}<\frac{n}{2}}\!\!\!\!\!\chi_{a,{\cal S}}(\texttt{q})\bigg{)}\ .$ (5.10)
Each term in the sum over ${\cal S}$ has multiplicity $(2{\cal S}+1)$ as a result of the $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ symmetry of the lattice model. Also, for every state with given $s=s_{a}$ (4.17) there exists another one with $s=-s_{a}$ which yields the same contribution to the partition function, except for the case when $s=\pm\frac{{\rm i}n}{2}$, where they are identified as the same state. This explains why the functions $\chi_{a,{\cal S}}(\texttt{q})$ come with a factor of two except the one with $a=-{\cal S}$ (recall that $s_{a}={\rm i}(\tfrac{n}{2}-{\cal S}-a)$ and hence $s_{a}=\frac{{\rm i}n}{2}$ for $a=-{\cal S}$).
@@
$Z^{({\rm cont})}=\sum_{{\cal S}\geq 0}(2{\cal S}+1)\,\int_{-\infty}^{\infty}\mbox{d}s\ \sum_{\texttt{d}\geq 0}\rho^{(\texttt{d})}_{p}(s)\,\texttt{q}^{\frac{s^{2}}{n}+\frac{p^{2}}{n+2}-\frac{1}{12}+\texttt{d}}\,.$ (5.11)
Here $\rho^{(\texttt{d})}_{p}(s)$ is the density of states defined in formulae (4.5) and (4.6), while recall that $2p=2{\cal S}+1-(n+2)$. Notice that $Z^{({\rm cont})}$ becomes singular as $L\to\infty$:
@@
$Z^{({\rm sing})}=\sqrt{\frac{n}{2\tau}}\ \frac{\log\big{(}2^{\frac{n+2}{n}}\,L/L_{0}\big{)}}{\pi\,\texttt{q}^{\frac{1}{24}}\prod_{m=1}^{\infty}(1-\texttt{q}^{m})}\ \sum^{\infty}_{\mathcal{S}=0}\,\,(2\mathcal{S}+1)\ \texttt{q}^{-\frac{1}{24}+\frac{p^{2}}{n+2}}\ \frac{1-\texttt{q}^{2\mathcal{S}+1}}{\prod_{m=1}^{\infty}(1-\texttt{q}^{m})}\,.$ (5.13)
@@
$\sum^{\infty}_{\mathcal{S}=0}\,\,(2\mathcal{S}+1)\ \texttt{q}^{-\frac{1}{24}+\frac{p^{2}}{n+2}}\ \frac{1-\texttt{q}^{2\mathcal{S}+1}}{\prod_{m=1}^{\infty}(1-\texttt{q}^{m})}\,,$ (5.14)
@@ 6 Discussion
In this work the universal behaviour of the staggered six-vertex model with $U_{q}\big{(}\mathfrak{sl}(2)\big{)}$ invariant boundary conditions imposed was considered. We focused on the so-called self-dual case in the critical regime (1.2). The problem was reformulated in the Hamiltonian picture, where a central rôle belongs to $\mathbb{H}$ (1), which commutes with the transfer-matrix of the vertex model. The study of the $1/L$ corrections to the $L\to\infty$ behaviour of the energy of $\mathbb{H}$ for the low energy states allowed us to extract the spectrum of scaling dimensions of the statistical system. Our treatment, involving the use of novel numerical and analytic techniques, represents an advance on the type of analysis of the scaling limit of integrable lattice systems with open boundary conditions that typically exists in the literature.
The numerical construction of the RG trajectories was achieved via the method of the $Q$ operator. A key rôle was played by the formula (2.20) for the matrix elements of $\mathbb{Q}(\zeta)$, valid for a one parameter family of open boundary conditions in the sector $S^{z}=0$. This is an original result of our work. The advantage of (2.20), as opposed to the expressions for $\mathbb{Q}(\zeta)$ appearing in the literature [28, 29, 30, 31], is that it contains no infinite sums; works literally for any (generic) complex values of $q$ and the boundary parameter $\epsilon$; and can be programmed efficiently on the computer. We believe that, in view of the important applications of the $Q$ operator, it may be worthwhile to extend (2.20) to the other sectors of the Hilbert space with $S^{z}\neq 0$ and the case of more general open boundary conditions.
@@
As was already observed in refs. [22, 23], the set of scaling dimensions of the statistical system possesses a continuous component, labeled by the quantum number $s=\operatorname*{slim}_{L\to\infty}b(L)\in\mathbb{R}$ with $b(L)$ from eq. (3.6) [23]. One of the results of this work is the explicit formula for the density of states $\rho^{(\texttt{d})}(s)$ (4.5), (4.6), which characterizes the continuous spectrum. In addition, we studied the RG trajectories $\{|\Psi_{L}\rangle\}$, where $s$ becomes a pure imaginary number in the scaling limit. Building on the analysis of [23], the discrete set $\Sigma\equiv\Sigma^{+}\cup\Sigma^{-}$ (4.19) of all admissible values of pure imaginary $s$ was found. We also determined the dimension of the linear span of states occuring in the scaling limit of $|\Psi_{L}\rangle$ with given $s\in\Sigma$ and conformal dimensions $\Delta$ (3.13).
Our work includes a full characterization of the linear space ${\cal H}$ appearing in the scaling limit of the space of low energy states of the lattice system. To describe it, ${\cal H}$ was decomposed into a direct sum of the ‘continuous spectrum’ ${\cal H}^{({\rm cont})}$ and the ‘discrete spectrum’ ${\cal H}^{({\rm disc})}$. The former, when expressed in terms of finite dimensional spaces, involves a direct integral over $s$, see eqs. (4.23), (4.10), (4.11), while the latter contains a direct sum (4.18). We explained how the graded linear spaces $\oplus_{\texttt{d}=0}^{\infty}{\cal H}_{p,s}^{({\rm cont},\texttt{d})}$ and $\oplus_{\texttt{d}=0}^{\infty}{\cal H}_{p,s}^{({\rm disc},\texttt{d})}$ are irreps. of the ${\cal W}_{\infty}$ algebra – the algebra of extended conformal symmetry of the model.
@@ Acknowledgments
The authors are grateful to M. Flohr and V. Schomerus for their interest in the work and for valuable discussions. In addition, we would like to thank S. Ribault for including an enlightening discussion about the CFT interpretation of our results in his referee report. GK would also like to thank V. V. Bazhanov, S. L. Lukyanov and J. Teschner for useful and stimulating interactions.
@@ Appendix A The asymptotic coefficients $\mathfrak{C}_{p,s}^{(\pm)}$
Here we provide a closed form expression for the coefficients $\mathfrak{C}_{p,s}^{(\pm)}(\boldsymbol{w})$ that were obtained in the work [45]. Among other things, they enter into the quantization condition (3.41).
@@
$\mathfrak{C}^{(\pm)}_{p,s}(\boldsymbol{w})=\mathfrak{C}_{p,s}^{(0,\pm)}\,\check{\mathfrak{C}}^{(\pm)}_{p,s}(\boldsymbol{w})\,,$ (A.1)
@@
[…]
ˇ$\check{\mathfrak{C}}^{(\pm)}_{p,s}(\boldsymbol{w})=\frac{(\mp 1)^{\texttt{d}}\ \det\big{(}w_{a}^{b-1}\,{U}^{(\pm)}_{a}(b)\big{)}}{\prod_{a=1}^{\texttt{d}}w_{a}\ \prod_{b>a}(w_{b}-w_{a})\ \prod_{a=1}^{\texttt{d}}\big{(}2p+2a-1\pm 2{\rm i}s\big{)}}$ (A.3)
@@
${U}^{(\pm)}_{a}(D)$ $=$ $(D-1)^{2}-\bigg{(}2p+2+n\mp 2w_{a}+\sum_{b\not=a}^{\texttt{d}}\frac{4w_{a}}{w_{a}-w_{b}}\bigg{)}\ (D-1)$
@@
$+$ $\bigg{(}\sum_{b\not=a}^{\texttt{d}}\frac{2w_{a}}{w_{a}-w_{b}}\,\bigg{)}^{2}+\,\big{(}2\,(2p+1+n\mp 2w_{a})-n\,\big{)}\,\sum_{b\not=a}^{\texttt{d}}\frac{w_{a}}{w_{a}-w_{b}}\ .$